# Learning from Emotions, Demographic Information and Implicit User Feedback in Task-Oriented Document-Grounded Dialogues

### v1

@@ Appendix B Dataset Features
In this section, we provide additional details on the demographic information and the error and user reaction types used to create FEDI.
@@
We distinguish 12 different language styles, including Their Age and Job, Standard, Colloquial, Formal, Gutter, Polite, Informal, Regional Dialect, Social Dialect, Jargon, Slang, Age, five demogrphic cohorts, including Boomers (born between 1952 and 1962), Generation X (born between 1962 and 1977), Millenials (born between 1977 and 1992), Generation Z (born between 1992 and 2007), and Generation Alpha (born between 2007 and 2016), and an extensive variety of occupations (overall 1,155), sampled from The Gazette [^17] Available in GitHub (last accessed on 31 July 2023)., including among others jobs from the fields of science and technology, education, arts and entertainment, healthcare, or manufacturing. For each dialogue, we randomly sample a new value for each characteristics and apply simple plausibility checks, e.g., a person from Generation A can only be a pupil.
Error and User Reaction Types
To generate implicit user feedback, we use the error and user response type taxonomy proposed by Petrak et al. (2023). For generation errors in system utterances they define the following nine error types as relevant for task-oriented document-grounded dialogues:
  • Ignore Question — This error occurs when the system fails to address the user’s question. Instead of providing a relevant response or clarification, the system disregards the user’s input.
@@
  • Ignore Expectation — This error happens when the system fails to fulfill the user’s expectations in terms of understanding and addressing their needs or requests accurately withing the context of the task.
  • Attribute Error — If the system fails to correctly extract or understand the necessary slots or attributes from the user’s utterance, this is called an attribute error.
@@
  • Unclear Intention — This error is characterized by the robot’s failure to accurately address the user’s intended objective.
  • Lack of Sociality — If a system’s response doesn’t adhere to social conventions, fails to include basic greetings, or exhibit toxic and disrespectful behavior or language, this is called Lack of Sociality.
They also define an error type for common-sense errors, but found them to be rare in task-oriented document-grounded dialogues. For this reason, we do not consider this error type in our work.
For user reactions in response to such generation errors, they propose the following taxonomy:
@@
According to the authors, Repeat or Rephrase, Make Aware with Correction, and Ask for Clarification contain implicit user feedback.
@@ Appendix C Prompts for Dialogue Generation and Annotation
In this section, we provide more details on the steps and prompts used for generating and annotating the dialogues of FEDI. Additionally added source data is highlighted in blue in the figures below.
@@
As described in Section 4, we require GPT-3.5-Turbo to return all results in a predefined JSON scheme, which depends on the prompt, i.e., dialogue generation or annotation, and ensures that the returned values contain all required fields and is processable without human intervention. If the values returned by GPT-3.5 do not adhere to the required scheme, we drop the whole dialogue. Figure 6 shows an example for the annotation of emotions.
@@
We append these json schemes at the end of the prompts before sending them to GPT-3.5-Turbo. We basically provide the required fields and example values, and instruct it to return only utf-8 encoded characters and escape quotation marks (so that we can treat it as a string in Python). Please refer to our GitHub repository for all prompts and their json schemes1.
@@
For this, we provide GPT-3.5-Turbo with the demographic information, the role of the starting actor, and the task description. We require the model to use this information to generate a background story and to use this as an additional source for dialogue generation. We also instruct the model to return the utterance-level annotations for intents in this step.
@@
The generation of feedback dialogues requires feedback scenarios as an additional source, which we generate in a separate step before dialogue generation using GPT-3.5-Turbo. We generate three feedback scenarios using the same promt for each new dialogue. Figure 8 shows the instruction.
@@
We provide the model with the task description and a list of three error and user reaction types. We then instruct the model to combine them uniquely and to generate three non-exclusive feedback scenarios that together form a story in the context of the task description. We also require the model to return a description and reasoning for each feedback scenario. We then use the returned feedback scenarios for the generation of feedback dialogues, which is shown in Figure 9.
Figure 9: Instruction for generating erroneous dialogues.
The instruction is longer and more detailed than the one used for generating the feedback-free dialogues (Figure 7). For example, it explicitely describes how to process error situations. Another difference is the length limitation. While feedback-free dialogues are restricted to 13 turns, we require feedback dialogues to have at least 13 turns. In practice, the length of the feedback dialogues is similar to the length of the feedback-free dialogues, but we observed that feedback dialogues are likely to be cut off without this requirement.
Continual Learning Dialogues
To create dialogues for continual learning, we create three additional versions of each feedback dialogue, each resolving one of the feedback scenarios from the original dialogue (until all feedback scenarios are solved). For this, we regenerate the system utterance with generation error. Figure 10 shows the instruction.
Figure 10: Instruction to regenerate the system utterance to replace the one with the generation error.
The instruction includes the dialogue up to the erroneous system utterance, the name of the task and the respective document (in the case of question answering). We mask the erroneous system utterance using <mask>, since we found this to produce replacements more coherent to the dialogue history. After replacing the erroneous system utterance, we regenerate its slot values. We also take advantage of the three-utterance-span constraint (see Section 4.2) and remove the following two utterances to ensure that the dialogue flow is not corrupted. This way, the conversation continues with the next regular user user utterance.
We experimented with different approaches for fixing errors in generated dialogues:
  • Using the implicit user feedback and the task description and instruct GPT-3.5-Turbo to rewrite the whole dialogue.
  • Providing GPT-3.5-Turbo with the whole dialogue and only instruct it to rewrite the affected turn.
  • Using the respective feedback scenario as additional input to regenerated the erroneous system utterance.
However, these approaches mostly resulted in inconsistent dialogues and off-topic, unnatural or incorrect system utterances. During prompt engineering, we found that the feedback itself is negligible for resolving feedback scenarios using our prompt from Figure 10, since it includes the original task description and requires the new system utterance to be generated in a friendly and polite manner.
@@
Figure 11 shows our instruction used for generating slot values.
@@
For this, we provide GPT-3.5-Turbo with the complete dialogue and distinguish between slots for each role (person and system). The slots to be annotated are provided in lists (including example values). We also instruct the model to just use sequences from the dialogue as slot values (to avoid hallucinated slot values).
@@
Figure 12 shows the instruction for emotion generation. We generate emotions just based on the dialogue context. We do not provide additional information, such as examples for each of the emotion. However, we additionally provide the number of utterances in the dialogue and those related to the user.
@@
Appendix D Studies With Human Annotators
In this work, we conducted several studies that required human annotators: (1) the test data collection for FEDI, (2) the curation and assessment of the generated annotations and (3) the human evaluation of our trained models. While the human evaluation was conducted by one of the authors and an intern from our research group (who participated during his working hours), the test data collection and curation study was conducted with external participants who were hired for this purpose.
Application Criteria and Hiring Procedure
For participation, we required a formal application. Our criteria were as follows:
@@ Appendix D Studies With Human Annotators
We considered a background in NLP, interest in conversational AI and experience in data annotation as a plus. We did not restrict the job advertisment to our university. Also, we did not take gender into consideration. We asked all applicants that fulfilled those criteria to take part in a recruitment test, in which we asked them to collect and annotate dialogues in a self-chat manner, given a random task description from our work. We then assessed and ranked their results based on (1) time needed for one dialogue, (2) annotation completness, (3) number of turns per dialogue, (4) avg. utterance length.
Overall, we received 11 applications, that all fulfilled our criteria. Eight of them passed the recruitment test and were hired for an hourly salary of 12,95\$. While all participants took part in the test data collection only two were involved in the data curation study.
Test Data Collection
For test data collection, we ransomly assigned participants to groups of two to collect the dialogues in one hour sessions dedicated to one task. For each task, we provided the task description, including slots with examples and four persona profiles (combinations of demographic information) and background stories as inspiration. However, we encouraged them to think about own persona profiles and background stories. For emotions, we provided lists of available options. For question answering, we also provided a set of at least 20 question-document pairs per topic.
For data collection, we use a self-developed web-based platform that allows to collect dialogues between two humans or between a human and a language generation model. Figure 13 shows the user interface.
@@
[…]
@@ Appendix E FEDI– Additional Analysis
In this section, we provide additional analysis about the composition of FEDI. Overall, FEDI consists of 8,852 dialogues, divided into 1,988 feedback-free and 6,864 feedback dialogues. Table 9 shows the distribution of dialogues in the dataset. Test for the feedback-free dialogues refers to the human-human collected test data.
@@
Version 1 Version 2 Version 3 Version 4
Train Dev Test Train Dev Train Dev Train Dev Train Dev
@@
Language styles are almost equally weighted. In terms of occupations, the figure shows that jobs from the categories of business administration, service, industrial and manufacturing, and pupil largely outweigh the other categories, which makes sense in the context of the tasks and topics represented in FEDI  [^18] The original list did not provide categories. We generated them using GPT-3.5-Turbo.. Overall, we observe 693 unique job titles in FEDI. The figures do not show the distribution of names. We found 1,496 different names in the dialogues. 638 (42%) are unique, and 712 (47.59%) occur two to three times. The remaining 146 names occur four or more times throughout the entire dataset.
@@
The chart in Figure 15 show the distribution of emotions in the dialogues of FEDI (including error and test dialogues).
@@
With 40.5%, Neutral is the most common emotion, followed by Curiosity (27.5%). Frustration and Confusion are relatively rare. We observe them mostly in the feedback dialogues. Others refers to emotions that are represented $\leq 5\%$, including Anger, Disgust, Fear, Surprise, and Stress.
@@
Overall, we generated 4,714 feedback scenarios that are included in the 1,716 feedback dialogues of Version 1. Figure 16 shows the distribution of error and user reaction types.
Figure 16: Distribution of error and user reaction types in the feedback dialogues of FEDI.
Given that most of the dialogues are about question answering (Table 9), it is not surprising that Ignore Question is the most frequent error type. Table 10 shows the ten most commonly observed error and user reaction type combinations.
@@ 10 Lack of Sociality
Table 10: The table shows the most common error and user reaction type combinations included in FEDI.
@@
Ignore Question and Ignore Request are two of the most frequent error types. While we observe the first one more common in question answering dialogues, the second one is more common in the other tasks. For both we observe that Ignore and Continue is the most frequent user reaction type, followed by Repeat or Rephrase. Unclear Intention is an error type mostly observed in parcel shipping, top up prepaid SIM card, and access control. The most frequently observed user reaction to this is Ask for Clarification. Based on absolute numbers, Factually Incorrect is the rarest error type. It’s mostly observed in question answering and in combination with Make Aware With Correction.
Appendix F Additional Details

### v2

@@ Appendix B Dataset Features
In this section, we provide additional details on the demographic information and the error and user feedback types used to create FEDI.
@@
We distinguish 12 different language styles, including Their Age and Job, Standard, Colloquial, Formal, Gutter, Polite, Informal, Regional Dialect, Social Dialect, Jargon, Slang, and Age. For age ranges, we consider five demographic cohorts, including Boomers (born between 1952 and 1962), Generation X (born between 1962 and 1977), Millenials (born between 1977 and 1992), Generation Z (born between 1992 and 2007), and Generation Alpha (born between 2007 and 2016). For occupations, we use a list of 1,155 job titles sampled from The Gazette [^18] Available in GitHub (last accessed on 16 January 2024)., including among others jobs from the fields of science and technology, education, arts and entertainment, healthcare, or manufacturing. As a source for the names, we use the list of the 2,000 most popular American baby names in 2010191919Published by babymed.com (last accessed 12 February 2024).. For each dialogue, we randomly sample a new value for each characteristic and apply simple plausibility checks, e.g., a person from Generation Alpha can only be a pupil.
Error and User Feedback Types
To generate generation errors and implicit user feedback, we use the error and user feedback type taxonomies proposed by Petrak et al. (2023). For generation errors in system utterances they define the following nine error types as relevant for task-oriented and document-grounded dialogues:
  • Ignore Question — This error occurs when the system fails to address a user’s question. Instead of providing a relevant response or clarification, the system disregards their input.
@@
  • Ignore Expectation — This error happens when the system fails to fulfill the user’s expectation in terms of understanding and addressing their needs within the context of the task.
  • Attribute Error — If the system fails to correctly extract or understand the necessary slots or attributes from a user’s utterance, this is called an attribute error.
@@
  • Unclear Intention — This error is characterized by the system’s failure to accurately address a user’s intended objective.
  • Lack of Sociality — If a system’s response doesn’t adhere to social conventions, fails to include basic greetings, or exhibit toxic and disrespectful behavior or language, this is referred to as a lack of sociality.
They also define an error type for common sense errors, but found them rare in task-oriented and document-grounded dialogues. For this reason, we do not consider this error type in our work.
For user feedback in response to generation errors, they propose the following taxonomy:
@@ Appendix C Prompts for Dialogue Generation and Annotation
Prompt engineering played a major role in this work. The instructions used to generate the dialogues and annotations were continuously improved in an iterative process to generate valid data within the given parameters. This section only focuses on the final instructions used in this work. Additionally added source data is highlighted in blue in the figures below.
@@
As described in Section 4, we require GPT-3.5 to return all results in a predefined JSON scheme, which depends on the prompt, i.e., dialogue generation or annotation, and ensures that the returned values contain all required fields and is processable without human intervention. If the values returned do not adhere to the required scheme, we drop the whole dialogue. Figure 6 shows an example for the annotation of emotions.
@@
We append these json schemes at the end of the prompts. We basically provide the required fields and example values, and instruct the model to return only utf-8 encoded characters and escape quotation marks (so that we can treat it as a string in Python). Please refer to our GitHub repository for all prompts and their json schemes1.
@@
We provide GPT-3.5 with the demographic information, the role of the starting actor, and the task description. We require the model to use this information to generate a background story and to use this as an additional source for dialogue generation. We also instruct the model to return the utterance-level annotations for intents in this step.
@@
Figure 8 shows the instruction for the generation of feedback scenarios, which are required as an additional source for feedback dialogues.
@@
For each feedback dialogue, we generate three feedback scenarios using the same prompt in a separate step before dialogue generation. Figure 9 shows the instruction for the generation of feedback dialogues.
Figure 9: Instruction for generating feedback dialogues.
The instruction is longer and more detailed than the one used for generating the feedback-free dialogues (Figure 7). For example, it explicitely describes how to process feedback scenarios. Another difference is the length limitation. While feedback-free dialogues are restricted to 13 turns, we require feedback dialogues to have at least 13 turns. In practice, the length of the feedback dialogues is similar to the length of the feedback-free dialogues, but we observed that feedback dialogues are likely to be cut off without this requirement. We consider the generated dialogue as Version 1.
Resolving Feedback Scenarios
For each feedback dialogue (Version 1), we generate three additional versions of the same dialogue, each resolving one of the feedback scenarios. For this, we experimented with different ideas:
  • Using the implicit user feedback and the task description and instruct GPT-3.5 to rewrite the whole dialogue.
  • Providing GPT-3.5 with the whole dialogue and only instruct it to rewrite the affected turn.
  • Using the respective feedback scenario as additional input to regenerate the affected system utterance.
They all resulted in inconsistent dialogues and off-topic or unnatural system utterances. We found that using the dialogue context up to the affected system utterance, masking and regenerating this utterance (in a friendly and polite manner), leads to the best matching and most coherent replacements. Figure 10 shows the instruction.
Figure 10: Instruction for regenerating the system utterance to replace the one with the generation error.
It includes the dialogue context, the name of the task and the document if the task is question answering. Although GPT-3.5 has a long context length, we found that including the full task descriptions was distracting rather than improving the replacements. This means that the model can only use internal knowledge and information from the dialogue context for generating the replacements, which sometimes had a negative impact on the completeness of the slot annotations, e.g., for parcel shipping and topping up a prepaid SIM card (see Section 5).
After replacing the affected system utterance, we regenerate its slot values. We remove the following two utterances to ensure the dialogue flow is not corrupted (since they directly refer to the generation error). The conversation then continues with the next regular user utterance. Figure 13 shows an example dialogue from FEDI to illustrate this procedure.
@@
Figure 11 shows our instruction for generating the slot annotations.
@@
For this, we provide GPT-3.5 with the complete dialogue and distinguish between slots for each role (person and system). The slots to be annotated are provided in lists (including example values). We also instruct the model to just use sequences from the dialogue as slot values (to avoid hallucinated slot values).
@@
Figure 12 shows the instruction for emotion generation.
@@
We generate emotions just based on the dialogue context. We do not provide additional information, such as examples. However, we additionally provide the number of utterances in the dialogue and those related to the user.
Figure 13: Example dialogue from FEDI for illustrating our approach for resolving feedback scenarios. In each version, we keep the previous part of the dialogue, regenerate the affected system utterance and drop the following two utterances (the user feedback and the system utterance which addresses the user feedback), since they are directly related to the generation error.
Appendix D Test Data Collection and Curation Study
We hired student assistants for our test data collection and curation study. In this section, we want to provide more insights into the application criteria, hiring procedure, and data collection.
D.1 Application Criteria and Hiring Procedure
To participate, we required a formal application. Our criteria were as follows:
@@ D.1 Application Criteria and Hiring Procedure
We considered a background in NLP, interest in conversational AI and experience in data annotation as a plus. We did not restrict the job advertisement to our university. Also, we did not consider gender. We asked all applicants who fulfilled those criteria to participate in a recruitment test, in which we asked them to collect and annotate dialogues in a self-chat manner, given a task description from our work. We then assessed and ranked their results based on (1) time needed for one dialogue, (2) annotation completeness, (3) number of turns per dialogue, (4) avg. utterance length.
Overall, we received 11 applications that fulfilled our criteria. Eight passed the recruitment test and were hired for an hourly salary of 12,95\$. While all participated in the test data collection only two were involved in the data curation study.
D.2 Test Data Collection
The test data for FEDI was collected by eight computer science students in overall 136 paid working hours. We randomly assigned participants to groups of two to collect the dialogues in one hour sessions dedicated to one task. For each task, we provided the task description, including slots with examples and four persona profiles (combinations of demographic information) and background stories as inspiration. However, we encouraged them to think about own persona profiles and background stories. For user emotions, we provided them with a list of available options. For question answering, we provided them with the question-document pairs extracted from the POSTE Italiane data (Section A).
For data collection, we used a self-developed web-based platform that allows to collect and annotate dialogues between two humans. Figure 14 shows the user interface.
@@ D.2 Test Data Collection
Each message is annotated with the respective intent (orange or green, depending on the role). Slot annotations are highlighted in yellow, with the slot type as superscript. User emotion annotations are colored purple. For Question Answering, the chatpane also allows attaching a document to a message (a text file).
D.3 Data Curation Study
For the curation of the generated data, the procedure was different for feedback-free and feedback dialogues. For feedback-free dialogues, we asked the annotators to assess and correct (add/modify/delete) the generated slot and intent annotations per utterance, and their completeness on dialogue level (with respect to the task description). We assigned the annotators to the tasks and asked them to work through the corresponding dialogues provided in INCEpTION Klie et al. (2018). Figure 15 shows the user interface for intent and slot annotation curation.
Figure 15: User interface for intent and slot annotation curation in INCEpTION. It’s a parcel shipping dialogue and the annotation for country of destination (Germany in line eight) is misplaced, because this slot should be provided by the user, who has already mentioned it in line five.
For feedback dialogues, we asked the annotators to assess and correct the annotations for implemented feedback scenarios, i.e., the annotation for error type in the affected system utterance and the user feedback type in the subsequent user utterance. In addition to the information available in the user interface, we provided the annotators with the task descriptions (Appendix A). For feedback dialogues, we also provided them with the definitions of error and user feedback types (Appendix B).
D.4 Dialogue Collection: Human vs. LLM
In our human test data collection, eight students collected 326 test dialogues in 136 paid working hours. With an hourly salary of 12.95\$, this adds up to a cost of 1,761.20\$ (not including additional costs, such as for supervision). Generating and annotating 8,526 dialogues using GPT-3.5 cost 75.73\$, including API calls for prompt engineering and debugging. On average, collecting and annotating a human-human dialogue cost 5.40\$. Using GPT-3.5, it is 0.009\$. Based on this, collecting and annotating dialogues with human participants is rather uneconomic and inefficient. However, with 175B parameters, GPT-3.5 is an extremely large model. Without access to such a model, this might be different. In a preliminary study, we used Llama-30B Touvron et al. (2023a) for dialogue generation and annotation. We asked a student assistant from our lab to assess the results. They constantly rated the Llama-30B dialogues lower in terms of naturalness, coherence, engagement, task coverage, i.e., how close is the generated dialogue to the task description, and (turn) length (see Table 9).
Model Naturalness Coherence Engagement Task Coverage Length
GPT-3.5-Turbo 4.40 4.92 1.0 4.68 7,12
LLaMA-30B 3.12 3.52 0.8 3.52 3,24
Table 9: Result of our analysis comparing dialogues generated by GPT-3.5 and Llama-30B. Except for Engagement and Length, all measurments are based on a Likert scale from 1 (lowest rating) to 5 (highest rating).
We suspect that this is rather due to the differences in model size and context window. While GPT-3.5 has a context window of 4k tokens, Llama-30B has a context window of only 2k tokens. However, regardless of the model used, LLM-generated data oftentimes suffers from various kinds of hallucinations Zhang et al. (2023b); Ji et al. (2023), which makes data curation with humans inevitable. In our data curation study (Section 6), we learned that this is not only much easier for humans, they are also much more efficient in curating annotated dialogues than collecting and annotating them from scratch. For example, collecting and annotating one dialogue takes on average ten minutes and requires two humans. For GPT-3.5 it is only 90 seconds. Curating an annotated dialogue took on average four minutes and did not require a partner.
@@ Appendix E FEDI– Additional Analysis
In this section, we provide additional analysis about the composition of FEDI. Overall, FEDI consists of 8,852 dialogues, 1,988 feedback-free and 6,864 feedback dialogues. Table 10 shows the distribution of dialogues in the dataset. Test refers to the human-human collected test data.
@@
Train Dev Test Version 1 Version 2 Version 3 Version 4 Dev
@@
Language styles are almost equally weighted. For occupations, the figure shows that jobs from the categories of business administration, service, industrial and manufacturing, and pupil largely outweigh the other categories, which makes sense in the context of the tasks and topics represented in FEDI  [^20] The original list did not provide categories. We generated them using GPT-3.5.. Overall, we observe 693 unique job titles in FEDI. The figures do not show the distribution of names. We found 1,496 different names in the dialogues. 638 (42%) are unique, and 712 (47.59%) occur two to three times. The remaining 146 names occur four or more times in the entire dataset.
@@
The chart in Figure 17 shows the distribution of emotions in the dialogues of FEDI.
@@
With 40.5%, Neutral is the most common emotion, followed by Curiosity (27.5%). Frustration and Confusion are relatively rare. We observe them mostly in the feedback dialogues. Other refers to emotions that are represented $\leq 5\%$, including Anger, Disgust, Fear, Surprise, and Stress.
@@
Overall, we generated 4,714 feedback scenarios that are included in the feedback dialogues of Version 1. Figure 18 shows the distribution of generation error and user feedback types.
Figure 18: Distribution of generation error and user feedback types in the feedback dialogues of FEDI.
Given that most of the dialogues are about question answering (Table 10), it is not surprising that Ignore Question is the most frequent error type. Table 11 shows the ten most commonly observed error and user feedback type combinations.
@@ 10 Lack of Sociality
Table 11: The table shows the most common error and user feedback type combinations included in FEDI.
@@
Ignore Question and Ignore Request are two of the most frequent error types. While we observe the first one more common in question answering dialogues, the second one is more common in the other tasks. For both we observe that Ignore and Continue is the most frequent user feedback type, followed by Repeat or Rephrase. Unclear Intention is an error type mostly observed in parcel shipping, topping up a prepaid SIM card, and access control. The most frequently observed user feedback to this is Ask for Clarification. Based on absolute numbers, Factually Incorrect is the rarest error type. It is mostly observed in question answering and in combination with Make Aware With Correction.
E.1 Curation Study – Results Analysis
To further investigate the quality of the generated annotations, we provide the human curation results in this section. Table 12 investigates the intent and slot annotations of the feedback-free dialogues before and after curation.
Task Intents Slots
Non-Curated Curated Non-Curated Curated
Parcel
Shipping
0.63 1.0 0.48 0.62
Top Up
SIM Card
0.71 1.0 0.61 0.91
Access
Control
0.95 1.0 0.38 0.60
Question
Answering
0.96 1.0 0.60 1.0
Table 12: The table compares the completeness with respect to intent and slot values before and after human curation.
As said in Section 6, the slot and intent annotations in the parcel shipping dialogues were most affected by human curation. The intent annotations for these dialogues are now complete. The completeness of the slot values was increased by 0.14 to 0.62. For access control, the situation is similar. The ratio of slot annotations in the Top Up SIM Card dialogues was increased by 0.29, and the question answering dialogues are now fully annotated.
Figure 19: Illustration of the distribution of emotions before and after human curation.
Figure 19 shows the changes in emotion annotations in the curated feedback-free dialogues. The ratio of each emotion has increased slightly, primarily due to the correction of unknown emotion annotations by the human curators, i.e. emotions that were not included in our taxonomy. The figure does not include the Neutral emotion which is still the most dominant emotion in the dataset (approx. 46% in the curated data).
Figure 20: Illustration of the distribution of user feedback types to generation error types before human curation.
Figure 20 shows the distribution of user feedback to generation error types before human curation. It shows a strong tendency towards Ignore and Continue in the case of Ignore Expectation, Ignore Request and Ignore Question errors. Figure 21 shows that many of these annotations were changed by the human curators to Make Aware With Correction and Repeat or Rephrase. The significant changes concerning Attribute Error, Ignore Question, Conversationality and Topic Transition Error show that the error type annotations were corrected particularly frequently in these cases.
Figure 21: Illustration of the distribution of user feedback types to generation error types after human curation.
The total number of error annotations itself has only changed slightly. We assume that the examples used for annotation were not distinct enough to support GPT-3.5 in interpreting the definitions for user feedback and generation errors (see Figure 2; for feedback scenario generation, we provided the model with definitions for error types and user feedback and examples). Especially in the case of e.g. Ignore Question and Ignore Request (see Appendix B for the definitions), it depends very much on specific details (e.g., punctuation marks).
Appendix F Experimental Details and Additional Results